# Understanding and Characterizing Obfuscated Funds Transfers in Ethereum Smart Contracts

### v1

@@ 4.1. System Overview
T1. Multi-step address generation F1. Number of steps in the addr generation
T2. Complex string operations F2. Number of string operations in the addr generation
@@
T3. External contract calls F3. Presence of an external contract call in the addr/value generation
@@
T4. Control‐flow complexity F4. Max branch/loop nesting depth in address/value derivation.
T5. Camouflage instructions F5. Transfer-related instruction ratio (TIR)
@@
T6. Replicated transfer logic F6. Inter-function code similarity of transfer-containing functions
T7. Irrelevant log events F7. Relevance between the semantics of the log event and the transfer operation
4.2. Definition and Extraction of Obfuscation Features
@@
For each of the aforementioned seven obfuscation strategies, we define a corresponding obfuscation feature, as summarized in Table 1.
@@
F1. Number of steps in address generation. This feature represents the number of steps required to obtain addr in a transfer operation. Specifically, starting from each CALL operation in the contract’s SSA IR, we perform backward dataflow analysis on the variable address to trace its generation process. Each arithmetic operation, hash function invocation, bitwise manipulation, and external call is counted as one distinct step. Finally, we consolidate linear operations within each basic block to avoid overcounting trivial operations and count the number of steps for address generation as a numerical feature F1.
@@
F2. Number of string operations. String operations (e.g., concatenation, hashing, slicing, and encoding) contribute to obfuscation. To quantify the level of complexity in string operations involved in the address generation process, we again analyze the data flow of the variable address to count all instructions involving string manipulations, including both built-in string operations and hash operations. We then count the number of string operations for address generation as a numerical feature F2.
@@
F3. Presence of external call. To determine if the generations of addr and value involve any external call, we inspect the dataflow of both and check if there exists any external call (i.e., (CALL, DELEGATECALL, or STATICCALL)). If so, we set a binary feature F3 to TRUE. Otherwise, it is set to FALSE.
@@
F4. Height of branch tree. We use the height of the branch tree to represent the complexity of the control flow associated with the transfer operation by analyzing conditional branch structures (JUMPI instructions). Starting from the identified transfer operation, we trace backward to traverse all conditional branches. The height of the branch tree is calculated as the maximum nesting depth traversed.
[…]
(2) $\text{TIR}=\frac{|U|}{|\text{T}|},$
@@
Specifically, $U$ is the union of critical transfer instructions $C$ and state modifications $S$. $C$ is obtained from the function’s program dependence graph (PDG) by considering all instructions directly related to the transfer logic (e.g., CALL, parameter calculations, state validation). $S$ is from parsing the function to detect SSTORE operations and retaining only those updating parameters related to transfers (e.g., address, value). Consequently, a low TIR implies that a large portion of the code in a transfer-residing function does not affect how the transfer operation is performed, indicating significant obfuscation designed to dilute and hide the contract’s true transfer logic.
@@
F6. Transfer operation similarity. This feature quantifies whether a contract replicates the same transfer logic across multiple functions. To calculate this, we construct a PDG, including information from both control and data dependencies, for each function containing a transfer operation. Then, we embed the nodes of each PDG using a Word2Vec model and further employ a Relational Graph Convolutional Network (R-GCN) (Schlichtkrull et al., 2018) to aggregate the node embeddings and edge relations into a single vector representation for each function’s subgraph. Finally, we compute pairwise cosine similarity between these vector representations across all transfer-containing functions and use this similarity score as a numerical feature F6. A high similarity score indicates that there exist two transfer-containing functions that exhibit highly similar transfer logic.
F7. Relevance of log events. To detect misleading and irrelevant log events emitted during transfer operations, we identify emit operations within two hops in the CFG from the node representing the transfer operation. If an emit operation is identified, we employ the Foundry fuzzing framework (Mota et al., 2023) to execute the corresponding function and capture the emitted logs at runtime. After that, we perform an automated semantic analysis using GPT-4o (Achiam et al., 2023) to determine whether the log messages are relevant to the transfer operation. If the relevance score is low (e.g., referencing a different recipient or transaction type), we set the feature to FALSE. Otherwise, the feature is set to TRUE.
Overall, we believe the above seven bytecode-level features can allow us to effectively identify obfuscated transfer-related functions in smart contracts.
4.3. Obfuscation Z-score Model
@@
Upon extracting the seven obfuscation features, we calculate an Obfuscation Z-score as a quantitative metric to represent the degree of obfuscation applied to transfer operations. In particular, we calculate the cumulative distance between each feature’s standard deviation and mean, and further compute a weighted sum to obtain the obfuscation Z-score for each contract:
(3) $Z_{\text{score}}=\sum_{i=1}^{7}\frac{x_{i}-\mu_{i}}{\sigma_{i}},$
where $x_{i}$ is the $i$-th feature value for a given contract, and $\mu_{i}$ and $\sigma_{i}$ are the mean and standard deviation of the corresponding feature across the entire sample set.
@@
This Z-score represents the cumulative distance between the values of all seven obfuscation features and their means in standard deviations, indicating the degree of obfuscation applied to the transfer operation for a given smart contract.
4.4. Evaluation of ObfProbe
@@
We evaluate the performance of ObfProbe on a dataset of 453 Ponzi scam contracts. Table 2 summarizes the data sources and our manual classification.
Table 2. Real-world Ponzi Scam Dataset
Data Source CRGB (Liang et al., 2024) 137
SourceP (Lu et al., 2024) 316
@@
To evaluate the effectiveness of ObfProbe, we compile these 453 contracts to obtain their bytecode and apply ObfProbe to get the values of the seven obfuscation features as well as the Z-scores. We manually confirm that the values of the seven obfuscation features are all correct, indicating 100% accuracy of our bytecode analysis over the Ponzi dataset. Furthermore, we examine the distribution of the two groups’ Z-scores, which is presented in Figure 2. As shown, the mean Z-score of the obfuscated group is significantly higher than that of the non-obfuscated group, and its standard deviation is also larger. This indicates that obfuscated contracts exhibit greater diversity and concealment.
In addition, we conduct a more detailed statistical investigation, shown in Table 3. The results show that the difference between the Z-score distributions of the two groups is statistically significant (t = -6.172, $p\approx 0$), confirming that ObfProbe can effectively distinguish obfuscated contracts from non-obfuscated contracts by using Z-scores.
Figure 2. Z-score Distribution
Table 3. Z-score Statistics
@@
Welch’s t-test t = -6.172, p = 0.000000
@@
In summary, the evaluation result shows that ObfProbe can accurately extract the pre-defined obfuscation features and calculate the Z-scores, which can effectively differentiate obfuscated and non-obfuscated smart contracts.
5. Real-World Study
@@
In this section, we aim to answer RQ2 to study the prevalence of obfuscation techniques in smart contracts.
5.1. Prevalence Study
To conduct the study, we collect 1,042,923 smart contract bytecodes deployed on the Ethereum mainnet spanning from Jun 2022 to Oct 2024, and then leverage ObfProbe to extract the seven obfuscation features from these contracts as well as to calculate their Z-scores. Next, we examine the top 0.3% most heavily obfuscated contracts to identify key hidden behaviors and their intrinsic obfuscation mechanisms. Finally, we present three case studies, namely MEV bot contracts, scam contracts, and extremely centralized contracts, to illustrate representative obfuscation strategies.
Figure 3. The Z-score Distribution on the Ethereum Mainnet. The dashed red line marks the top 0.3% cutoff.
@@ 5.1. Prevalence Study
Distribution of Z-Score. Figure 3 shows that the distribution of the aggregate obfuscation score z_total is strongly right‐skewed. Most contracts fall in the range $3\lesssim Z\lesssim 8$, with a peak near $Z\approx 6$. For $Z>10$, the bar heights decline steadily, forming a long tail that extends to the highest observed scores. This visual pattern indicates that while most contracts exhibit moderate obfuscation, a small subset reaches much higher levels.
Table 4. Zscore Statistics
@@
Table 4 summarizes the distribution of the aggregate obfuscation score (z_total) across 1,042,923 contracts. To examine the prevalence of obfuscated contracts, we use a Z-score threshold of 4.637 derived from real-world data in Table 2 and represents the 95% CI upper bound, calculated as
[…]
Table 5. Prevalence of Obfuscated Smart Contracts
@@
Above Threshold (>4.637) 739,763 70.93 %
Below Threshold (<4.637) 303,160 29.07 %
Applying this cutoff, Table 5 reveals that 70.93% of on-chain contracts deploy obfuscation at or above what would be considered a “normal” level. More than two-thirds of deployed contracts exceed the non-obfuscated CI ceiling, demonstrating that obfuscation has become a routine practice in smart contract development. This pervasive adoption underscores the need for more robust analysis tools and transparency mechanisms to manage obfuscation in the Ethereum ecosystem.
5.2. Detailed Analysis
[…]
To this end, we sort these contracts by the total amount of influx funds and examine each contract from the top. Our study uncovers that all of them are indeed equipped with obfuscation techniques, and 153 contracts pose very high risks by hiding four potentially malicious and suspicious behaviors via heavy obfuscation, jeopardizing  100M USD worth of Ethers. We present case studies for the four malicious behaviors hidden under obfuscation and elaborate on their typical hidden strategies.
5.3. Case I: MEV Bots

### v2

@@ IV-A System Overview
T1. Multi-step address generation F1. Number of steps in the addr generation.
T2. Complex string operations F2. Number of string operations in the addr generation.
@@
T3. External contract calls F3. Presence of an external contract call in the addr/value derivation.
@@
T4. Control-flow complexity F4. Max branch/loop nesting depth along addr/value derivation.
T5. Camouflage instructions F5. Transfer-related Instruction Ratio (TIR).
@@
T6. Replicated transfer logic F6. Inter-function similarity among transfer-containing functions.
T7. Irrelevant log events F7. Semantic relevance between log events and the transfer operation.
IV-B Definition and Extraction of Obfuscation Features
@@
For each of the aforementioned seven obfuscation strategies, we define a corresponding obfuscation feature that can be extracted by ObfProbe, as summarized in Table I.
@@
F1. Number of steps in address generation. This numerical feature represents the number of steps required to obtain addr in a transfer operation. Starting from each CALL operation in the contract’s SSA IR, we perform backward dataflow analysis on the parameter address to trace its generation process. Each arithmetic operation, hash function invocation, bitwise manipulation, and external call is counted as one distinct step. Finally, we consolidate linear operations within each basic block to avoid overcounting trivial operations, and we treat the number of steps for address generation as a numerical feature F1.
@@
F2. Number of string operations. String operations (e.g., concatenation, hashing, slicing, and encoding) contribute to obfuscation. To quantify the complexity of string operations in the address generation process, we reanalyze the data flow of the parameter addr to count all instructions that involve string manipulation, including both built-in string operations and hash operations. We then count the number of string operations for address generation as a numerical feature F2.
@@
F3. Presence of an external call. To determine if the generations of addr and value involve any external call, we inspect the dataflow of both and check if there exists any external call (i.e.,CALL), we set a binary feature F3 to TRUE, otherwise, to FALSE.
@@
F4. Height of the branch tree. We use the height of the branch tree as a numerical feature to represent the complexity of the control flow associated with the transfer operation by analyzing conditional branch structures (JUMPI instructions). Starting from the transfer operation, we backtrace to traverse all conditional branches. The height of the branch tree is calculated as the maximum nesting depth traversed.
F5. Transfer-related instruction ratio (TIR). This feature quantifies the ratio of effective instructions contributing to a transfer operation in the residing function. To this end, we define Transfer-related Instruction Ratio (TIR) as: $\text{TIR}=\frac{|U|}{|T|}$, where $|U|$ is the number of effective instructions contributed to transfers and critical state updates, and $|T|$ is the total number of instructions in the transfer-residing function. Specifically, $U$ is computed as the union of critical transfer instructions $C$ and state modifications $S$.
@@
We obtain $C$ from the function’s program dependence graph (PDG) by considering all instructions directly related to the transfer logic (e.g., CALL, parameter calculations, state validation). $S$ is from parsing the function to detect SSTORE operations and retaining only those updating parameters related to transfers (e.g., address, value). Consequently, a low TIR indicates that a large portion of the code in a transfer-resident function does not affect how the transfer operation is performed, suggesting significant obfuscation designed to dilute and hide the contract’s true transfer logic.
@@
F6. Transfer operation similarity. This feature quantifies whether a contract replicates the same transfer logic across multiple functions. To calculate this, we construct a PDG that includes information from both control and data dependencies for each function containing a transfer operation. Then, we embed the nodes of each PDG using a Word2Vec model and further employ a Relational Graph Convolutional Network (R-GCN) [51] to aggregate the node embeddings and edge relations into a single vector representation for each function’s subgraph. Finally, we compute pairwise cosine similarities between these vector representations across all transfer-containing functions and use these scores as numerical feature F6. A larger similarity score indicates that two transfer-containing functions implement more similar transfer logic. We treat this feature as a continuous indicator rather than a binary decision, allowing for some noise in the functions.
F7. Relevance of log events. F7 captures whether log events emitted along transfer logic are semantically unrelated to the transfer itself (i.e., noisy or misleading logs). We decode log topics/data when ABI signatures are available and semantically classify each log as transfer-related or not using an LLM-assisted classifier. Our design includes a rollback mechanism that resolves ambiguous cases (i.e., those with low LLM confidence scores) via manual annotation. In a 100-sample pilot, the LLM achieves 100% accuracy on this binary relevance task, so we use it to support semantic labeling. We set F7 to TRUE if irrelevant logs co-occur with transfer operations, otherwise FALSE.
IV-C Obfuscation Z-score Model
@@
Upon extracting the seven obfuscation features, we customize the standard Z-score [16] and compute an Obfuscation Z-score as a quantitative metric for the degree of obfuscation applied to transfer operations. In particular, we calculate the cumulative distance between each feature’s standard deviation and mean, and further compute the sum of standardized features to obtain the obfuscation Z-score for each contract:
$Z_{\text{score}}=\sum_{i=1}^{7}\frac{x_{i}-\mu_{i}}{\sigma_{i}},$ (2)
@@
where $x_{i}$ is the $i$-th feature value for a given contract, and $\mu_{i}$ and $\sigma_{i}$ are the mean and standard deviation of the corresponding feature across the entire sample set. This Z-score represents the cumulative distance between the values of all seven features and their means, expressed in standard deviations, indicating the degree of obfuscation applied to the fund transfer operations.
Why Z-score. Our seven features reside on heterogeneous scales (e.g., counts, ratios, graph-similarity scores), so summing raw values would be disproportionately influenced by the feature with the largest numeric range. Standardizing each feature using its corpus-level mean and standard deviation renders them unitless and comparable, ensuring that the aggregate score reflects joint extremeness rather than scale artifacts. The resulting Z-score quantifies how many standard deviations a contract’s combined behavior deviates from the corpus mean in the current snapshot, enabling corpus-agnostic tail selection (e.g., “top three-sigma”). Operationally, re-estimating $(\mu_{i},\sigma_{i})$ for each snapshot allows the score to adapt to ecosystem or compiler drift without manual, per-feature threshold tuning. In sum, Z-score normalization promotes fairness across features, interpretability across datasets, and stability over time.
IV-D Evaluation of ObfProbe
@@
We evaluate the performance of ObfProbe on a dataset of 453 Ponzi-scam contracts, focusing on its effectiveness. Table III summarizes the data sources and our manual classification results.
TABLE II: Real-world Ponzi Scam Dataset
Category Item Count
Data source CRGB [34] 137
SourceP [38] 316
@@
TABLE III: Z-score statistics on the labeled dataset.
@@
Welch’s t-test: t=6.172t=-6.172, $p<10^{-6}$
To evaluate the effectiveness of ObfProbe, we compile the 453 contracts to obtain their bytecode and apply ObfProbe to obtain the values of the seven obfuscation features, as well as the Z-scores. We manually check the extracted F1–F7 signals on these 453 contracts and find no extraction errors (feature-correctness check, $N=453$), which is different from detection accuracy. Furthermore, we examine the distribution of the Z-scores for the two groups (obfuscated vs. non-obfuscated), which is presented in Figure 2. As shown, the mean Z-score of the obfuscated group is significantly higher than that of the non-obfuscated group, and its standard deviation is also larger. This indicates that obfuscated contracts exhibit greater diversity and concealment. Additionally, we conduct a more detailed statistical investigation, as shown in Table III. The results show that the difference between the Z-score distributions of the two groups is statistically significant (t = -6.172, $p\approx 0$), confirming that ObfProbe can effectively distinguish obfuscated contracts from non-obfuscated contracts by using Z-scores.
Table IV summarizes standard detection metrics (i.e., precision, recall, ROC-AUC, and PR-AUC) for the proposed Z-score rule and a logistic-regression baseline trained on features F1–F7. Precision is the proportion of contracts predicted as obfuscated that are truly obfuscated, whereas recall is the proportion of obfuscated contracts correctly identified. The Z-score cutoff achieves perfect recall (1.0000) but low precision (0.2921), indicating a high false-positive rate; in contrast, logistic regression attains substantially higher precision (0.6130) at the cost of lower recall, reflecting the expected false-positive/false-negative trade-off on the same labeled set. Because the labels are imbalanced (92 obfuscated vs. 361 non-obfuscated), PR-AUC is more informative than accuracy, and moderate F1 scores are therefore expected in this setting. For the Z-score method, ROC-AUC and PR-AUC are computed by treating the raw Z-scores as continuous confidence values, while precision/recall/F1 are reported at the fixed threshold $Z\geq 4.637$. This threshold is derived from the labeled non-obfuscated subset of the Ponzi dataset as the upper bound of the 95% confidence interval and then applied uniformly. Importantly, it is not tuned to optimize evaluation metrics. All logistic regression results are computed from out-of-fold predictions obtained under 5-fold cross-validation.
[…]
@@
In summary, the evaluation results show that ObfProbe can accurately extract the predefined obfuscation features. In addition, it also shows that our Z-score representation model can effectively differentiate obfuscated from non-obfuscated smart contracts.
TABLE IV: Detection metrics on the Ponzi dataset.
Method Precision Recall F1 ROC-AUC PR-AUC
Z-score cutoff ($\geq$ 4.637) 0.2921 1.0000 0.4521 0.8463 0.6627
LogReg (CV=5, 7 signals) 0.6130 0.4550 0.5186 0.7683 0.5888
TABLE V: Drop-column feature importance (Top-3).
Feature $\Delta$ PR-AUC $\Delta$ F1
F5 0.0744 0.1028
F3 0.0261 0.0291
F2 0.0222 0.0305
Figure 2: Z-score distribution on Ponzi scam dataset.
Figure 3: Z-score distribution on mainnet (red dashed = top 3000 cutoff).
Static Analysis Efficiency. The first pass yields SSA outputs for 1,042,923 contracts (93.5474% first-run success) under a 20s timeout. The first-pass runtime averages 19.66s with a median of 19.9s. Failed contracts are diagnosed and rerun with the timeout removed; recovered outputs are merged back, so the prevalence analysis includes these rerun samples, indicating a stable pipeline. In the uncapped rerun, the average runtime is about 45s (median 41s), and large/complex bytecodes often fall in the 20–80s range, so the runtime is dominated by bytecode size and complexity.
V Smart Contract Obfuscations in the Wild
@@
In this section, we aim to answer RQ2 by studying the prevalence of obfuscation techniques in real world smart contracts.
V-A Prevalence Study
We collect 1,042,923 unique smart contract bytecodes that were active on Ethereum mainnet between Jun 2022 and Oct 2024 (i.e., at least one on-chain interaction in this window), regardless of their original deployment dates.
@@ V-A Prevalence Study
Distribution of Z-Score. Figure 3 shows that the distribution of the obfuscation score (Z-score is strongly right‐skewed. Most contracts fall in the range $3\lesssim Z\lesssim 8$, with a peak near $Z\approx 6$. For $Z>10$, the bar heights decline steadily, forming a long tail that extends to the highest scores. This visual pattern indicates that while most smart contracts exhibit a moderate level of obfuscation, a small subset employs an extremely high level. In Table VII, we summarize the distribution of Z-score across the 1,042,923 contracts. To examine the prevalence of obfuscated contracts, we use a Z-score threshold of 4.637 derived from real-world data in Table III to represent the 95% confidence interval (CI) upper bound, calculated as
$\vskip-5.05942pt4.571\;+\;t_{0.975,360}\times\frac{0.641}{\sqrt{361}}\;\approx\;4.637.$ (3)
Applying this cutoff, Table VII reveals that 70.93% of on-chain contracts deploy obfuscation at or above what would be considered a “normal” level. More than two-thirds of deployed contracts exceed the non-obfuscated CI ceiling, indicating that obfuscation has become routine in smart contract development. This pervasive adoption underscores the need for more robust analysis tools and transparency mechanisms to manage obfuscation in the Ethereum ecosystem.
TABLE VI: Z-score statistics.
@@
TABLE VII: Prevalence of obfuscated smart contracts.
@@
Above threshold (>4.637>4.637) 739,763 70.93%
Below threshold (<4.637<4.637) 303,160 29.07%
These results highlight that the key distinction in practice is degree rather than mere presence: most contracts cluster in a moderate range, whereas the most atypical behavior concentrates in the extreme right tail. The fact that 70.93% of contracts exceed the non-obfuscated ceiling suggests that moderate obfuscation is commonplace, motivating prioritization of manual analysis on the top-3000 contracts with the highest Z-scores. This corpus-level prevalence provides a baseline for the remainder of our study: we focus on the extreme tail, where rare signals are concentrated, and auditing effort is expected to yield the highest return.
Prevalence of individual features. Table VIII reports, for each feature, its non-zero rate $P(F_{i}>0)$ and strength quantiles (median/P90/P99) over 1,042,923 contracts. F1, F4, and F5 occur in more than 70% of contracts, indicating that multi-step address derivation, control-flow complexity, and camouflage instruction ratio are widespread in real-world bytecode. Consequently, feature presence alone is not a reliable risk indicator and should be interpreted jointly with magnitude (e.g., P90/P99). The distributions are markedly heavy-tailed (e.g., F1 increases from a median of 4 to a P90 of 478), showing that extreme contracts amplify these primitives by orders of magnitude. By contrast, F3 appears in only 8.50% of contracts, suggesting that external-call obfuscation is less prevalent but potentially more discriminative when it occurs.
TABLE VIII: Feature prevalence in 1,042,923 contracts.
Feature P(F¿0) Median P90 P99
F1 72.91% 4.00 478.00 974.00
F2 41.20% 0.00 120.00 314.00
F3 8.50% 0.00 0.00 1.00
F4 73.67% 4.00 25.00 43.00
F5 73.82% 0.14 0.56 0.87
F6 58.79% 56.03 98.77 99.90
F7 39.79% 0.00 1.00 1.00
Prevalence of obfuscation patterns. Using the T1–T7 $\leftrightarrow$ F1–F7 mapping, Table IX reports pattern prevalence across the full corpus. T1, T4, and T5 are the most common (each $\approx$73%), indicating that these primitives constitute a broad baseline of real-world obfuscation. T2 and T6 are moderately prevalent (41.20% and 58.79%), whereas T3 is comparatively rare (8.50%), consistent with external-call obfuscation being less frequently deployed. As high-frequency base patterns, T1/T4/T5 are not, by themselves, strong indicators of maliciousness; in contrast, the rarer T3 pattern can provide a sharper signal in prioritization settings.
TABLE IX: Pattern prevalence in 1,042,923 contracts.
Pattern P_all
T1 72.91%
T2 41.20%
T3 8.50%
T4 73.67%
T5 73.82%
T6 58.79%
T7 39.79%
Decomposition of Z-score. Table X decomposes the overall Z-score into feature-level shares, computed as mean absolute standardized contributions. F3 accounts for the largest share (18.04%), followed by F7 (16.69%) and F6 (16.33%), indicating that these components contribute disproportionately to the aggregate score. F2 contributes the smallest share (10.90%). Overall, high Z-scores are driven primarily by strong, less ubiquitous signals rather than by the most common background patterns.
TABLE X: Z-score contribution share across the corpus.
Feature Share_all
F1 12.03%
F2 10.90%
F3 18.04%
F4 13.13%
F5 12.88%
F6 16.33%
F7 16.69%
Temporal trend (2022–2024). To quantify the trend for RQ2, we associate each contract with its creation timestamp and compute monthly aggregates from June 2022 through August 2024 (UTC). Figure 4 plots the monthly median Z-score (top) and the prevalence of each feature (bottom). The median Z-score remains largely stable around $\sim 5.0,$ while individual feature prevalences vary within bounded ranges, suggesting a stable aggregate baseline with shifts in the relative emphasis of specific techniques. Notably, the prevalence curves for F1/F4/F5 nearly overlap (differences <0.5<0.5 percentage points), consistent with these high-frequency primitives co-occurring as baseline obfuscation; consequently, their month-to-month movements are tightly coupled. Our dataset includes contract creations up to 2024-08; subsequent months contain no new deployments in this snapshot. These observations motivate the extreme-tail analysis in Section V-B.
Figure 4: Temporal trend of obfuscation signals (2022–2024): monthly median Z-score (top) and P(Fi¿0) for all features (bottom).
Answer to RQ2. Applying ObfProbe to real-world Ethereum smart contracts reveals that most real-world smart contracts exhibit a moderate level obfuscation, and there exists a small subset employing an extreme higher level of obfuscations;
V-B Analysis of the 3000 highly-obfuscated Contracts
To better understand the impact of obfuscation techniques used in real-world smart contracts, we select the top 3000 contracts in terms of Z-scores as highly suspicious targets for in-depth analysis. We focus our analysis on top 3000 contracts based on two considerations. First, in statistics, the rightmost 3000 samples (roughly 0.3% of the total) in a normal distribution typically indicate extreme outliers [66, 14] (i.e., values exceeding approximately three standard deviations from the mean). Second, focusing on the top 3,000 samples can significantly reduce our manual efforts while still adequately covering the typical values of the distribution. We hence limit our manual analysis and case studies to the top 3,000 contracts to strike a balance between feasibility and coverage.
Overview of the top 3000 contracts: Table XII presents an overview of the top 3000 contracts. It can be seen that the top 3000 contracts’ Z-score is heavy-tailed, evidenced by a median of 21.922, a minimum of 20.419, and a maximum of 46.264. In addition, all 3000 contracts (100%) have exhibited at least 2 obfuscation features, and 98.1% have exhibited at least 3 obfuscation features. Among the 3000 contracts, 46.0% have their source code published on Etherscan. We distinguish explicit obfuscation (verified source code visible to users) from covert obfuscation (bytecode-only or unverified). In the top-3000 tail, only 46.0% are explicit by this definition, leaving the majority covert. This coupling between deep obfuscation and code invisibility raises audit cost and motivates focusing manual efforts on the extreme tail. Table XII shows the distribution of the top 3000 contracts’ Z-score. It can be seen that the majority of the contracts’ Z-score fall into the range of 20–25 (88.1%). 11.1% of them fall into the range of 25 - 30. Less than 1% of them fall into the range of 30 - 50. In short, the top-3000 set is the actionable slice for auditing: stacked obfuscation and low source visibility make it the highest-yield target for further investigation.
TABLE XI: Top 3000 obfuscated contracts.
Metric Value (Share)
$z_{\text{score}}$ min / median / max 20.419 / 21.922 / 46.264
features $\geq 2$ 3,000 (100.0%)
features $\geq 3$ 2,943 (98.1%)
With verified source code 1,380 (46.0%)
TABLE XII: Distribution of $z_{\text{score}}$ in top 3000 contracts.
$z_{\text{score}}$ range Count (Share)
20–25 2,644 (88.1%)
25–30 335 (11.1%)
30–35 16 (0.5%)
35–40 2 (0.1%)
40–45 2 (0.1%)
45–50 1 (0.0%)
Top-3000 feature prevalence. Table XIII shows that the extreme tail is dominated by F1/F2/F4/F5/F6 (near-universal presence), while F3 exhibits strong enrichment (11.44x) and F7 remains present in 72.87% of top-3000 contracts. The near-universal coverage indicates that the extreme tail reflects stacked obfuscation rather than a single trick, and F3’s enrichment makes external-call hiding a key gate into the tail.
TABLE XIII: Top-3000 feature prevalence and enrichment.
Feature P_top3k Enrichment Median_top3k
F1 100.00% 1.37x 1199.00
F2 100.00% 2.43x 417.00
F3 97.23% 11.44x 1.00
F4 100.00% 1.36x 33.00
F5 100.00% 1.35x 0.11
F6 99.13% 1.69x 98.40
F7 72.87% 1.83x 1.00
Ablation on top-3000 stability. Removing F3 yields the largest overlap drop (40.43%), followed by F2 (27.37%) and F4 (19.37%), indicating these signals most strongly shape the extreme tail ranking. By contrast, F5/F6/F7 have much smaller drops; they act as background noise rather than decisive ranking drivers.
TABLE XIV: Top-3000 ranking stability under feature ablation.
Feature Overlap@3000 Drop
F1 81.20% 18.80%
F2 72.63% 27.37%
F3 59.57% 40.43%
F4 80.63% 19.37%
F5 96.23% 3.77%
F6 97.70% 2.30%
F7 96.33% 3.67%
TopK sensitivity. The enrichment trends remain stable across Top-1000/3000 (Table XV), indicating the extreme tail is a structural phenomenon rather than a threshold artifact.
TABLE XV: TopK sensitivity for selected features.
K F3 F7
1000 98.30% 72.00%
3000 97.23% 72.87%
A close look at the top 3000 contracts. To facilitate further analysis, we examine the top 3000 contracts ranked by Z-score and use total funds to prioritize risk assessment during manual inspection. Our manual investigation reveals that all contracts employ obfuscation techniques, with 463 contracts posing very high risks due to extensive obfuscation that conceals four potentially malicious and suspicious behaviors. Such 463 contracts can be divided into four categories: MEV bots, Ponzi schemes, fake decentralization, and extreme centralization. The total amount of funds absorbed by these contracts reaches approximately \$100 million USD in Ether. Table XVI summarizes the distribution of the 463 contracts among the four categories. In general, MEV bots account for the largest share (50.1%), followed by extreme centralization (33.0%). Ponzi and Fake Decentralization are less frequent but salient for risk analysis. Below, we conduct a detailed case study of the four contract types, discussing their typical obfuscation patterns, associated financial impacts, and the lifespan of their transaction activity.
TABLE XVI: Distribution of the contracts in the four case studies.
Case-study pattern Count Percentage
MEV bots 232 50.1%
Extreme Centralization 153 33.0%
Ponzi 48 10.4%
Fake Decentralization (renounce) 30 6.5%
Total instances (multi-label) 463 100%
V-C Case I: MEV Bots